Evaulate (-3) second power -17
A) 39
B) 24
C) 27
D)30

Answer by fractalier(6550)

(-3)^2 - 17 =
9 - 17 =
-8
Even if you meant -3^2 - 17, that answer is -9 - 17 = -26
There is a mistake somewhere...